OpenStudy (anonymous):

An airplane flew 3 hours with a 40 mph head wind. The return trip with a tail wind of the same speed took 2 hours. Find the speed of the plane in still air

OpenStudy (anonymous):

We can use the formula d = rt The first equation is d = (x + 40)3 What did you get for the second equation?

OpenStudy (anonymous):

We know that d=vt (d=distance, v=velocity (speed), and t=time) and also that the distance each way is exactly the same. Therefore we can set the Distances equal to each other: D = D Vt*time = Vh * time (Vp+40)2 = (Vp-40)3 2Vp+80 = 3Vp -120 200 = Vp So the speed of the plane in still air is 200mph.